OpenStudy (anonymous):

ok i have a little stupid question :-S what is the volume of z^2=X^2+y^2 with z belonging to [0;1] triple integrals :-S

OpenStudy (precal):

triple integrals are pretty smart material.......

OpenStudy (anonymous):

its not a stupid question at all :), find some solutions of similar problems on the internet so you can get some point how to calculate volume in R3 space

OpenStudy (anonymous):

done that, it's just that this one is kind of awkward, i think it should be 2pi/3 but from my calculation that is the only thing that isn't

OpenStudy (experimentx):

\[\frac 2 3 \pi\]

OpenStudy (experimentx):

woops!! z goes from 0 to 1, it's 1/3 pi

OpenStudy (experimentx):

easy way,  another way calculate this integral \[ \int_{-1}^{1} \int_{-\sqrt{1 - x^2}}^{\sqrt{1 - x^2}}\int_{0}^{\sqrt{x^2 + y^ 2}} dz dy dz\]